The following is ABC Inc.’s balance sheet (in thousands

Assets Cash Accounts receivable Inventory $ 20 90 90 Liabilities and Equity Accounts payable $ 30 Notes payable 90 Accruals L Also, sales equal $500, cost of goods sold equals $360, interest payments equal $62, taxes equal $56, and net income equals $22. The beginning retained earnings is $0, the market value of equity is equal to its book value, and the company pays no dividends.
Calculate Altman’s Z score for ABC, Inc. if ABC has a 50 percent dividend payout ratio and the market value of equity is equal to its book value. Recall the following: Net working capital
=
Current assets − Current liabilities
Current assets
=
Cash + Accounts receivable + Inve ntories
Current liabilities
=
Accounts payable + Accruals + Notes payable
EBIT
=
Revenues − Cost of goods sold − Depreciation
Taxes
=
(EBIT − Interest)(Tax rate)
Net income
=
EBIT − Interest − Taxes
Retained earnings
=
Net income(1 − Dividend payout ratio)
B) Should you approve ABC Inc.’s application to your bank for $500,000 for a capital expansion loan?
C)If ABC’s sales were $450,000, taxes were $16,000, and the market value of equity fell to one-quarter of its book value (assume cost of goods sold and interest are unchanged), how would that change ABC’s income statement? With the new data, does your credit decision change?
D)What are some of the shortcomings of using a discriminant function model to evaluate credit risk
Answer #1
Z-Score = 1.2A + 1.4B + 3.3C + 0.6D + 1.0E
Where:
A = working capital / total assets
B = retained earnings / total assets
C = earnings before interest and tax / total assets
D = market value of equity / total assets
E = sales / total assets
A = ((20+90+90)-(30+90+30))/700
A = 0.07143
B = 22/700
B = 0.03143
C = (22+56+62)/700
C = 0.20
D = 400 / 700
D = 0.57143
E = 500 / 700
E = 0.71429
Z-Score = 1.2A + 1.4B + 3.3C + 0.6D + 1.0E
Z-Score = (1.2*0.07143) + (1.4*0.03143) + (3.3*0.20) + (0.6*0.57143) + (1.0*0.71429)
Z-Score = 1.846866
I would reject the loan application as the z score is 1.846866 which is just above 1.8 and if the z score is below 1.8 means firm is heading toward bankruptcy. If score is above 3 then we would have accepted loan application

Part C)

Sales 450
COGS 360
Gross profit 90
Interest 62
Earning before taxes 28
Taxes 16
Net income 12
A = ((20+90+90)-(30+90+30))/700 0.071429
A = 0.07143
B = 12/700 0.017143
B = 0.01714
C = (62+28)/700 0.128571
C = 0.12857
D = (400/4) / 700 0.142857
D = 0..14286
E = 450 / 700 0.642857
E = 0.64286
Z-Score = 1.2A + 1.4B + 3.3C + 0.6D + 1.0E
Z-Score = (1.2*0.07143) + (1.4*0.01714) + (3.3*0.12857) + (0.6*0.14286) + (1.0*0.64286)
Z-Score = 1.26527

Under new sales Z score is 1.26527 which is below 1.8 so company is heading towards bankruptcy.